(9,5)and(10,7) find the slope of the line that goes through
Answer by stanbon(75887)

(9,5)and(10,7) find the slope of the line that goes through
---------------
slope = (7-5)/(10-9) = 2/1 = 2
==================
Cheers,
Stan H.
=============